40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 26 15:07:47 2016, Steve, Update, PEM, accelerometers moved to ETMX ETMX_floor_vs_optab.pngETMX_floor_vs_optabtop.png
    Reply  Wed Jan 27 08:17:38 2016, Steve, Update, PEM, ETMX floor vs table noise 6x
       Reply  Wed Jan 27 20:51:00 2016, rana, Update, PEM, ETMX floor vs table noise 
          Reply  Wed Jan 27 23:14:25 2016, gautam, Update, PEM, ETMX floor vs table noise X_end_ACC_Data.pngX_End_ACC_Data.xml.tar.gz
Message ID: 11950     Entry time: Tue Jan 26 15:07:47 2016     Reply to this: 11952
Author: Steve 
Type: Update 
Category: PEM 
Subject: accelerometers moved to ETMX 

Objective:  measure the noise floor on the optical table and the floor so we can decide if the table needs better anchoring before swapping in

                   the larger optical table

The accelerometrs labeled as MC1 ( just north east  of IOO chamber floor ) and MC2 ( north east leg of MC2 table floor ) were moved:

MC1 to the floor at the north west leg of optical table.

MC2 is in the north east corner of the optical table

Atm2 was taken after table leg bolts were tighed at 40 ft/lb

The spectrum looks similar to ETMY     (Guralp for below 3 Hz )  except  the Z direction 

ETMY  .

Attachment 1: ETMX_floor_vs_optab.png  90 kB  | Show | Hide all | Show all
Attachment 2: ETMX_floor_vs_optabtop.png  98 kB  Uploaded Tue Jan 26 16:22:13 2016  | Hide | Hide all | Show all
ETMX_floor_vs_optabtop.png
ELOG V3.1.3-